A tunnel is dug along a diameter of the earth.Find the force on a particle of mass m placed in the tunnel at a distance x from the centre.

Grade:11

1 Answers

Sarah Saha
8 years ago
   F=Gm1m2/rxr (where,m1 sthe mass of the earth & m2 is the mass of the particle & r is the distance between them)------------(1)
      Again,F=m2a=m2g---------------(2)
       Therefore from equation 1 & 2,we get:-
          g=Gm1/rxr (where r is the radius of the earth )
 g=G \frac {m_1}{r^2}=(6.67384 \times 10^{-11}) \frac{5.9722 \times 10^{24}}{(6.371 \times 10^6)^2}=9.8196 \mbox{m} \cdot \mbox{s}^{-2}